# From Compression to Expansion: A Layerwise Analysis of In-Context Learning

### v1

@@ Abstract
In-context learning (ICL) enables large language models (LLMs) to adapt to new tasks without weight updates by learning from demonstration sequences. While ICL shows strong empirical performance, its internal representational mechanisms are not yet well understood. In this work, we conduct a statistical geometric analysis of ICL representations to investigate how task-specific information is captured across layers. Our analysis reveals an intriguing phenomenon, which we term Layerwise Compression-Expansion: early layers progressively produce compact and discriminative representations that encode task information from the input demonstrations, while later layers expand these representations to incorporate the query and generate the prediction. This phenomenon is observed consistently across diverse tasks and a range of contemporary LLM architectures. We demonstrate that it has important implications for ICL performance—improving with model size and the number of demonstrations—and for robustness in the presence of noisy examples. To further understand the effect of the compact task representation, we propose a bias-variance decomposition and provide a theoretical analysis showing how attention mechanisms contribute to reducing both variance and bias, thereby enhancing performance as the number of demonstrations increases. Our findings reveal an intriguing layerwise dynamic in ICL, highlight how structured representations emerge within LLMs, and showcase that analyzing internal representations can facilitate a deeper understanding of model behavior.
@@ 1 Introduction
Recent research has advanced several theoretical perspectives on explaining why ICL works—viewing ICL as posterior inference [3], implicit meta-learning [4], internal optimization [5, 6] or mechanistic interpretations[7]. However, the underlying mechanism of how LLMs distinguish different tasks and use this information to guide their output remains unclear for ICL. To address this gap, we focus on the hierarchical feature learning across layers and formulate the following research question:
@@
To investigate how task-specific information evolves across layers, we conduct a statistical geometric analysis of ICL representations across multiple tasks. Specifically, we consider a set of $T$ ICL tasks, each associated with a distribution $\mathcal{D}_{t}$ for input-output pairs. Then for each task $t$, we randomly sample $K$ input-output pairs (also referred to as demonstrations) from $\mathcal{D}_{t}$, which are combined with a query to form an ICL instance. We construct multiple such instances per task following this procedure. To quantify how the model compresses task information in its internal representations, we examine two key aspects: (1) how instances from the same task are clustered together, and (2) how instances from different tasks are distinguished from each other. Our analysis reveals an intriguing phenomenon, which we term the Layerwise Compression-Expansion, summarized as follows:
Layerwise Compression-Expansion Phenomenon LLMs exhibiting ICL capabilities organize their layers into two parts with distinct behaviors: a compression part and an expansion part. The early layers, comprising the compression part, progressively produce compact and discriminative representations that capture task information from the input demonstrations. The later layers, forming the expansion part, apply these compact representations to the query to generate the output.
@@
Figure 1: Layer-wise compression to expansion in ICL representations. TDNV first decreases then increases from shallow to deep layers, splitting the model into compression and expansion stages. During the compression stage, task vector accuracy increases as task information is progressively extracted from demonstration pairs. During the expansion stage, early-exit accuracy increases as output information is progressively decoded based on the input query. Refer to Appendix F for detailed explanation of task vector and early-exit accuracy.
As shown in Figure 1, TDNV first decreases and then increases from shallow to deep layers, splitting the model into compression and expansion stages. To further support our hypothesis, we measure task vector accuracy [9] and early-exit accuracy [10, 11] across layers to track task-specific and output-specific information. Task vector accuracy measures zero-shot ICL performance when injecting intermediate layer hidden states extracted under ICL settings. Early-exit accuracy measures performance by directly applying the final-layer classifier to intermediate hidden states. During compression, task vector accuracy increases while early-exit accuracy remains low, indicating that the model is compressing representations to encode task information. During expansion, task vector accuracy decreases while early-exit accuracy rapidly increases, indicating that the model begins to incorporate query information and decode output-specific representations. As we show, this behavior has important implications for ICL performance and robustness.
@@
  • By analyzing the hidden representations of in-context learning, we conceptualize and extensively examine the Layerwise Compression-Expansion phenomenon. Our results show that this phenomenon is prevalent across different tasks and a range of contemporary large language models, including both transformer-based architectures and state space models.
@@
  • We show that noisy demonstrations result in less compressed representations and a corresponding drop in performance. However, the hidden representations remain distinguishable with a certain amount of noise, which helps explain the robustness of ICL. Moreover, we find that errors in early demonstrations can be suppressed by later examples, and that errors in later demonstrations lead to less compressed representations than those in early ones. This sheds light on the recency phenomenon [12, 13] and highlights the greater importance of later demonstrations.
@@ 2 Related Works
Layerwise Representations An intriguing line of research [18, 19, 20, 21, 22, 23] has empirically investigated the role of different layers in feature learning. These studies show that in image classification tasks, features in intermediate layers become increasingly linearly separable as the layers deepen. Specifically, Neural Collapse ($\mathcal{NC}$) properties emerge in intermediate layers, where the within-class variance decreases compared to the between-class variance as depth increases. This indicates that layerwise compression occurs monotonically with layer depth in these settings. However, our hypothesis reveals that in the ICL setting, decoder-only models’ layerwise representations exhibit dual encoding-decoding stages: shallow layers compress information while deep layers expand it. Furthermore, research by [24] shows that intermediate layers consistently outperform both shallow and final layers on downstream tasks. Our research focuses on how intermediate layers achieve maximum information compression in the ICL setting.
In-Context Learning Interpretability Numerous studies have focused on the working mechanisms of ICL [3, 4, 5, 25, 6, 7]. Xie et al. [3] propose that ICL emerges as an implicit form of Bayesian inference. In the realm of meta-learning, Chen et al. [4] introduce in-context tuning to predict target labels from concatenated sequences of task instructions. A significant line of research connects ICL with gradient descent optimization. Von Oswald et al. [5] demonstrate that Transformers trained on autoregressive tasks can emulate gradient descent in their forward pass. Dai et al. [25], Ahn et al. [6] compare standard gradient descent-based fine-tuning with ICL, revealing that transformer attention in ICL exhibits a dual form of gradient descent-based optimization. Olsson et al. [7] identify "induction heads"—specific attention heads that facilitate ICL by copying patterns from previous tokens. However, our work focuses on the layer-wise representational analysis of ICL.
@@ 3 Preliminaries
In this section, we first formally set up the layer-wise representations of in context learning in Section 3.1, followed by introducing the metrics for measuring within-task compression of features at each layer in Section 3.2.
@@ 3.1 Problem Setup
Layerwise ICL Representations For ICL task, we are given $(i)$ $K$ demonstrations, denoted as $\mathcal{S}_{K}=\{s_{1},s_{2},\dots,s_{K}\}$, where each demonstration $s_{k}=(x_{k}\rightarrow y_{k})$ consists of an input token $x_{k}$, a separator token $\rightarrow$", and a label token $y_{k}$; and $(ii)$ a query input $\mathcal{X}=(x_{q}\rightarrow)$. We refer to the demonstration-query pair $(\mathcal{S}_{K},\mathcal{X})$ as an ICL instance. An LLM $f$ performs ICL by processing the instance $(\mathcal{S}_{K},\mathcal{X})$ as a next-token prediction task. Let $\bm{Z}^{(\ell)}\in\mathbb{R}^{d\times p}$ denote the hidden representations at layer $\ell$ for the instance, where $p$ denotes the sequence length and $d$ represents the dimension of the hidden representation. The layerwise update with $f$ is performed as
@@ 3.2 Metric for Representation Compression
The study of feature compression and discrimination [30, 19, 31, 32, 33, 8, 20, 34] has recently gained significant attention in representation learning. Inspired by this line of work, we analyze how models compress task information in their internal representations by examining two key aspects.
@@
4 Layerwise Compression-Expansion Dynamic
@@
In this section, we examine the dynamics of layer-wise representation under the in-context learning setting, a phenomenon we termed Layerwise Compression-Expansion. The subsequent sections validate and explore this phenomenon in detail across various conditions. Specifically, Section 4.2 demonstrates that it occurs universally across different architectures. Next, we analyze key factors influencing this phenomenon, including model size (Section 4.3) and demonstration noise (Section 4.4).
4.1 Experimental Setup
To examine this phenomenon, we designed our experimental setup following Hendel et al. [9], covering commonly used architectures and in-context learning (ICL) tasks.
Models We evaluate the hypothesis on two distinct architectural families: ($i$) Decoder-only transformers, including Llama3 [35], Pythia [36], GPT-J [37] and Deepseek-coder [38]. ($ii$) State-space models, specifically Mamba [39]. Given that Deepseek-coder demonstrates stronger ICL performance on symbolic tasks, we use Deepseek-coder-7B[38] as our default model unless otherwise specified.
Tasks Consider two sets of tasks: (1) Letter-to-Letter, including 5 tasks (copy letter, next letter, to uppercase, previous letter and next 2 letter). For all tasks, the query is fixed and only demonstrations are different. (2) List-to-Element, including 5 tasks (list first, list last, list length, list first upper, list last upper). For each task, we sample $N=100$ instances, setting the default number of demonstrations to $K=15$. Detailed descriptions of all tasks are provided in Appendix A.
4.2 Prevalence of Layerwise Compression-Expansion
Layer-wise compression-expansion as a general mechanism of ICL. Our analysis shows that the Layerwise Compression-Expansion phenomenon is consistently observed across various LLMs with different architectures when performing. As shown in Figure 3, the TDNV metric consistently exhibits a U-shaped trend—first decreasing then increasing—across a range of models, including transformers (Llama3 [35], Pythia [36], GPT-J [37], and Deepseek [38]) and state space models (Mamba [39]). This demonstrates that information is progressively compressed across early layers to form a compact task representation, and then expanded in deeper layers to guide output generation. This phenomenon holds even in the absence of attention, as evidenced by Mamba, indicating that the mechanism is not specific to the transformer architecture.
@@ 4.2 Prevalence of Layerwise Compression-Expansion
Figure 2: Layerwise TDNV of different model architectures, including transformer and state space model.
@@
Figure 3: Layerwise TDNV during training process. The Layerwise Compression-Expansion phenomenon emerges and intensifies with training progress.
To further characterize this mechanism at the token level, we use saliency maps[40] computed via gradient-weighted attention. Additional details about the saliency map can be found in Appendix C. In a nutshell, these maps highlight how token interactions evolve across layers. From Figure 4, we observe that $(i)$ in shallow layers, the model focuses on token-to-label mappings within demonstrations, reflecting local compression, $(ii)$ intermediate layers shift focus toward the final token, indicating integration of task-level information, and $(iii)$ deep layers emphasize interactions between the query and final token, aligning with expansion and prediction. Thus, token-level interpretability also aligns with the layerwise compression-expansion trajectory.
The most compressed layer aligns with the optimal task vector layer. To find the optimal intermediate layer for task vector extraction, previous works [9] typically patch the vector at each layer and evaluate accuracy using a validation set. TDNV provides a more efficient way to identify the optimal layer for task vectors. As shown in Figure 1, at a certain intermediate layer $\hat{\ell}$, we observe three simultaneous changes: the TDNV shifts from decreasing to increasing, the task vector accuracy begins to decrease, and the early-exit accuracy starts to increase. Since the layer with minimum TDNV corresponds to the layer with maximum task vector accuracy, we can identify the optimal layer with just one inference using ^$\hat{\ell}=\operatorname*{\text{arg~{}min}}_{\ell\in[1,L]}\text{TDNV}^{(\ell)}$.
@@
Emerges of Layerwise Compression-Expansion during training. To verify that the Layerwise Compression-Expansion emerges only in trained models and not in randomly initialized ones, we evaluate on the Amber-6.7B models of LLM360 family [41], which provide intermediate checkpoints throughout the training process. As shown in Figure 3, models with random initialization exhibit flat TDNV values across all layers, indicating no structure of information compression. As training proceeds, the TDNV curve transitions into a distinct U-shape: early layers compress task information (decreasing TDNV), and deeper layers expand it (increasing TDNV). This suggests that the model progressively learns to structure its internal representations by first compressing task information and then expanding it to support query-specific prediction—a pattern that only emerges as a result of training.
@@ 4.3 Scaling Up Model size Leads to More Compression
To explore how model size influences information compression, we analyze Pythia models ranging from 14M to 410M parameters in terms of both layerwise TDNV (as shown in Figure 6) and performance (in Figure 6). We evaluate ICL performance from two perspectives: (1) the regular few-shot setting, referred to as ICL, and (2) the task-vector (TV) setting—i.e., zero-shot ICL using a task vector patched from the best-performing layer $\hat{\ell}$ identified under the few-shot setting—referred to as TV ICL. Higher accuracy in either setting indicates better performance. Additionally, we report zero-shot accuracy without any task-vector information, referred to as the baseline. We find that larger models tend to produce more compressed and discriminative representations in the middle layers, indicating a stronger ability to extract task information from the demonstrations, thereby achieving better performance in both ICL and task-vector ICL.
Figure 5: Layerwise TDNV of varying model size.
Figure 6: ICL Performance v.s. minimum TDNV of varying model size.
4.4 Compression-to-Expansion under Noisy Demonstrations
@@
Demonstrations in real-world scenarios are often noisy, with some input-output pairs failing to accurately represent the intended task. Despite this, ICL demonstrates notable robustness to such noisy demonstrations. As illustrated in Figure 16(a), performance remains largely unaffected even when the noise ratio reaches up to 40%, where the noise ratio is defined as the proportion of incorrect input-output pairs relative to the total number of pairs. To understand this robustness, we explore it through the lens of information compression.
Figure 7: ICL Performance under different noise ratios.
Figure 8: Layerwise TDNV under different noise ratios.
@@
We plot the layerwise TDNV under varying noise ratios in Figure 8 and highlight two key observations: ($i$) higher noise ratios consistently lead to increased TDNV across all layers, indicating that noisy demonstrations impair the model’s ability to compress and extract task-relevant information. In the extreme case of 100% noise—where inputs and labels are completely uncorrelated—the model receives no meaningful task signal, and the characteristic compression-to-expansion pattern disappears across layers. ($ii$) When the noise ratio remains below 40%, the minimum TDNV values stay below 1, indicating that within-task variance is still smaller than between-task distance. This allows task representations to remain distinguishable, resulting in minimal performance degradation. This observation explains the robust performance at noise ratios below 40%, as shown in Figure 16(a). However, beyond 40% noise, task representations become increasingly entangled, causing performance to decline rapidly.
The position of noisy demonstration affects compression
Figure 9: ICL Performance and minimum TDNV of last separation token.
@@
Unlike conventional machine learning, the order of demonstrations has a significant impact on model performance in ICL [42, 43, 44]. As illustrated in Figure 9, perturbing demonstrations that appear later in the sequence causes a larger performance drop and higher TDNV values. These perturbations result in less compressed task vectors, indicating that later demonstrations play a more crucial role than earlier ones in helping the model extract task information.
To investigate why perturbing later demonstrations results in higher TDNV at the final separation token, we conduct a more fine-grained analysis by computing the layerwise TDNV for each separation token across all demonstrations. For each demonstration’s separator, we calculate the layerwise TDNV and organize them into a grid structure, referred to as the grid CDNV. As shown in Figure 10, perturbing a demonstration at a given position most significantly increases the TDNV of the next demonstration. However, this increase gradually diminishes as more correct demonstrations are appended. This pattern suggests that the negative impact of early errors can be partially mitigated by subsequent correct examples.

### v2

# From Compression to Expression: A Layerwise Analysis of In-Context Learning

@@ Abstract
In-context learning (ICL) enables large language models (LLMs) to adapt to new tasks without weight updates by learning from demonstration sequences. While ICL shows strong empirical performance, its internal representational mechanisms are not yet well understood. In this work, we conduct a statistical geometric analysis of ICL representations to investigate how task-specific information is captured across layers. Our analysis reveals an intriguing phenomenon, which we term Layerwise Compression-Expression: early layers progressively produce compact and discriminative representations that encode task information from the input demonstrations, while later layers express these representations to incorporate the query and generate the prediction. This phenomenon is observed consistently across diverse tasks and a range of contemporary LLM architectures. We demonstrate that it has important implications for ICL performance—improving with model size and the number of demonstrations—and for robustness in the presence of noisy examples. To further understand the effect of the compact task representation, we propose a bias-variance decomposition and provide a theoretical analysis showing how attention mechanisms contribute to reducing both variance and bias, thereby enhancing performance as the number of demonstrations increases. Our findings reveal an intriguing layerwise dynamic in ICL, highlight how structured representations emerge within LLMs, and showcase that analyzing internal representations can facilitate a deeper understanding of model behavior.
@@ 1 Introduction
Recent research has advanced several theoretical perspectives on explaining why ICL works—viewing ICL as posterior inference [3], implicit meta-learning [4], internal optimization [5] or mechanistic interpretations[6]. However, the underlying mechanism of how LLMs distinguish different tasks and use this information to guide their output remains unclear for ICL. To address this gap, we focus on the hierarchical feature learning across layers and formulate the following research question:
@@
To investigate how task-specific information evolves across layers, we conduct a statistical geometric analysis of ICL representations across multiple tasks. Specifically, we consider a set of $T$ ICL tasks, each associated with a distribution $\mathcal{D}_{t}$ for input-output pairs. Then for each task $t$, we randomly sample $K$ input-output pairs (also referred to as demonstrations) from $\mathcal{D}_{t}$, which are combined with a query to form an ICL instance. We construct multiple such instances per task following this procedure. To quantify how the model compresses task information in its internal representations, we examine two key aspects: (1) how instances from the same task are clustered together, and (2) how instances from different tasks are distinguished from each other. Our analysis reveals an intriguing phenomenon, which we term the Layerwise Compression-Expression, summarized as:
Layerwise Compression-Expression Phenomenon LLMs exhibiting ICL capabilities organize their layers into two parts with distinct behaviors: a compression part and an expression part. The early layers, comprising the compression part, progressively produce compact and discriminative representations that capture task information from the input demonstrations. The later layers, forming the expression part, apply these compact representations to the query to generate the output.
@@
Figure 1: Layer-wise compression to expression in ICL representations. TDNV first decreases then increases from shallow to deep layers, splitting the model into compression and expression stages. During the compression stage, task vector accuracy increases as task information is progressively extracted from demonstration pairs. During the expression stage, early-exit accuracy increases as output information is progressively decoded based on the input query. Refer to Appendix˜G for detailed explanation of task vector and early-exit accuracy.
As shown in Figure˜1, TDNV first decreases and then increases from shallow to deep layers, splitting the model into compression and expression stages. To further support our hypothesis, we measure task vector accuracy [8] and early-exit accuracy [9, 10] across layers to track task-specific and output-specific information. Task vector accuracy measures zero-shot ICL performance when injecting intermediate layer hidden states extracted under ICL settings. Early-exit accuracy measures performance by directly applying the final-layer classifier to intermediate hidden states. During compression, task vector accuracy increases while early-exit accuracy remains low, indicating that the model is compressing representations to encode task information. During expression, task vector accuracy decreases while early-exit accuracy rapidly increases, indicating that the model begins to incorporate query information and decode output-specific representations. As we show, this behavior has important implications for ICL performance and robustness.
@@
  • By analyzing the hidden representations of ICL, we conceptualize and extensively examine the Layerwise Compression-Expression phenomenon. Our results show that it is prevalent across model architectures (transformer and state-space models) and task domains (symbolic, language understanding and multimodality), and emerges during the training process.
@@
  • We show that noisy demonstrations result in less compressed representations and a corresponding drop in performance. However, the representations remain distinguishable with a certain amount of noise, which helps explain the robustness of ICL. Moreover, we find that errors in early demonstrations can be suppressed by later examples, and that errors in later demonstrations lead to less compressed representations than those in early ones. This highlights the recency effect [11, 12] and the key role of later demonstrations.
  • Motivated by our analysis, we propose task-vector contrastive fine-tuning method to further compress task vectors and reduce TDNV. Fine-tuning GPT-2 models on symbolic ICL tasks with this approach yields 20% improvement on average in task-vector accuracy over standard finetuning.
@@ 2 Related Works
Layerwise Representations. Prior works [17, 18, 19, 20, 21, 22] investigated the role of different layers in feature learning. They revealed that in classification task, intermediate layer features become increasingly linearly separable and exhibit Neural Collapse ($\mathcal{NC}$), indicating monotonic feature compression with depth. In contrast, we hypothesize that decoder-only ICL models follow a dual process where shallow layers compress information and deeper layers re-express it, and intermediate layers achieve maximal compression.
In-Context Learning Interpretability. Numerous studies have investigated the mechanisms underlying ICL [3, 4, 5, 23, 24, 6, 25], spanning perspectives of Bayesian inference, meta-learning, and optimization. Our work instead analyzes through layer-wise representations. In addition, Doimo et al. [26] examine the geometry of ICL representations by clustering intermediate features according to semantic subjects of input, whereas our findings differ by focusing on the underlying tasks induced by input–output pairs.
Task Representations. Various compact representations capture ICL tasks, including task vector [8], function vector [27], and state vector [28], which guide model behavior by injecting hidden states. Other works explore compositional and latent-space manipulation [29, 30]. Prior studies focus on single-task representations, whereas we provide a layer-wise geometric analysis of representations.
A more comprehensive discussion of the related works can be found in Appendix˜B.
@@ 3 Preliminaries
In this section, we first formally set up the layer-wise representations of in context learning in Section˜3.1, followed by introducing the metrics for measuring within-task compression of features at each layer in Section˜3.2.
@@ 3.1 Problem Setup
Layerwise ICL Representations. For ICL task, we are given $(i)$ $K$ demonstrations, denoted as $\mathcal{S}_{K}=\{s_{1},s_{2},\dots,s_{K}\}$, where each demonstration $s_{k}=(x_{k}\rightarrow y_{k})$ consists of an input token $x_{k}$, a separator token $\rightarrow"$, and a label token $y_{k}$; and $(ii)$ a query input $\mathcal{X}=(x_{q}\rightarrow)$. We refer to the demonstration-query pair $(\mathcal{S}_{K},\mathcal{X})$ as an ICL instance. An LLM $f$ performs ICL by processing the instance $(\mathcal{S}_{K},\mathcal{X})$ as a next-token prediction task. Let $\bm{Z}^{(\ell)}\in\mathbb{R}^{d\times p}$ denote the hidden representations at layer $\ell$ for the instance, where $p$ denotes the sequence length and $d$ represents the dimension of the hidden representation. The layerwise update with $f$ is performed as
@@ 3.2 Metric for Representation Compression
The study of feature compression and discrimination [31, 32, 33, 34, 7, 35] has recently gained significant attention in representation learning. Inspired by this line of work, we analyze how models compress task information in their internal representations by examining two key aspects.
@@
4 Layerwise Compression-Expression Dynamic
@@
In this section, we examine the dynamics of layer-wise representation under the ICL setting, a phenomenon we termed Layerwise Compression-Expression. The subsequent sections validate and explore this phenomenon in detail across various conditions. Specifically, Section˜4.1 demonstrates that it occurs universally across different architectures and tasks. Next, we analyze key factors influencing this phenomenon, including model size (Section˜4.2) and demonstration noise (Section˜4.3).
4.1 Prevalence of Layerwise Compression-Expression
To validate whether Layerwise Compression-Expression is a general mechanism of ICL, we evaluate it across different LLM model architectures and tasks. Unless otherwise specified, we use Deepseek-coder-7B [36] as our default model. For each task, we sample $N=100$ instances, setting the default number of demonstrations to $K=15$.
Universality across model architectures. Following Hendel et al. [8], we first evaluate the algorithmic domain, including 5 tasks (copy letter, next letter, to uppercase, previous letter and next 2 letter). Detailed descriptions of all tasks are provided in Appendix˜A. As shown in Figure˜3, the TDNV metric consistently exhibits a U-shaped trend—first decreasing then increasing—across two distinct architectural families: ($i$) Decoder-only transformers, including Llama3 [37], Pythia [38], GPT-J [39] and Deepseek-coder [36]. ($ii$) State-space models, specifically Mamba [40]. This phenomenon holds even in the absence of attention, as evidenced by Mamba, indicating that the mechanism is not specific to the transformer architecture.
@@ 4.1 Prevalence of Layerwise Compression-Expression
Figure 2: Layerwise TDNV of different model architectures, including decoder-only transformers and state-space models.
@@
Figure 3: Layerwise TDNV during training process. The phenomenon emerges and intensifies with training progress.
Universality across task domains. To evaluate the generality of the phenomenon beyond algorithmic settings, we examine three additional task categories: ($i$) Symbolic ICL. We adopt the linguistic, translation, and knowledge domains from Hendel et al. [8], where TDNV consistently exhibits a U-shaped trend (Figure˜6). ($ii$) Language Understanding ICL. Beyond only short phrases, we evaluated TDNV on a natural language dataset with longer sentences. Each sentence can be analyzed across multiple attributes: length, semantic polarity, tense, sentence type, subject person, and entity type. We adopt Llama3 8B [37] to predict the attribute label (e.g., positive or negative) for a query sentence based provided demonstrations with labels of a specific attribute (e.g., semantic polarity). As shown in Figure˜6, the TDNV also exhibits a U-shaped trend, with the most compressed representation shifting to a later layer (around layer 28), indicating that longer sentences require more layers for effective task compression. (iiiiii) Multimodality ICL. We further extend to a vision–language setting using a 2-D shape dataset, where each image contains a shape with four attributes (color, shape, size, texture). We adopt the Qwen-VL [41] model to predict the attribute label (e.g., red or green) for a query image based provided demonstrations with labels of a specific attribute (e.g., color). As shown in Figure˜6, the TDNV metric again exhibits a U-shaped curve. Across all settings, increasing the in-context length $K$ leads to more compact internal representations with lower TDNV values. Complete task specifications are provided in Appendix˜A.
Figure 4: Symbolic ICL.
Figure 5: Language Understanding ICL.
Figure 6: Multimodality ICL.
@@
Emergence during training. To verify that the phenomenon emerges only in trained models and not in randomly initialized ones, we evaluate on the Amber models of LLM360 family [42], which provide intermediate checkpoints throughout the training process. As shown in Figure˜3, models with random initialization exhibit flat TDNV values across all layers, indicating no structure of information compression. As training proceeds, the TDNV curve transitions into a distinct U-shape curve. This suggests the phenomenon only emerges as a result of training.
@@ 4.2 Scaling Up Model size Leads to More Compression
To explore how model size influences information compression, we analyze Pythia models ranging from 14M to 410M parameters in terms of both layerwise TDNV and performance (as shown in Figure˜8). We evaluate ICL performance from two perspectives: (1) the regular few-shot setting, referred to as ICL, and (2) the task-vector (TV) setting—i.e., zero-shot ICL using a task vector patched from the best-performing layer $\hat{\ell}$ identified under the few-shot setting—referred to as TV ICL. Higher accuracy in either setting indicates better performance. Additionally, we report zero-shot accuracy without any task-vector information, referred to as the baseline. We find that larger models tend to produce more compressed and discriminative representations in the middle layers, indicating a stronger ability to extract task information from the demonstrations, thereby achieving better performance in both ICL and task-vector ICL.
Figure 7: Effect of model size on layerwise TDNV and ICL performance.
Figure 8: Effect of noisy demonstrations on ICL performance and layerwise TDNV.
4.3 Compression-to-Expression under Noisy Demonstrations
@@
Demonstrations in real-world scenarios are often noisy, with some input-output pairs failing to accurately represent the intended task. Despite this, ICL demonstrates notable robustness to such noisy demonstrations. As illustrated in Figure˜8 left, performance remains largely unaffected even when the noise ratio reaches up to 40%, where the noise ratio is defined as the proportion of incorrect input-output pairs relative to the total number of pairs. To understand this robustness, we explore it through the lens of information compression.
@@
We plot the layerwise TDNV under varying noise ratios in Figure˜8 right and highlight two key observations: ($i$) higher noise ratios consistently lead to increased TDNV across all layers, indicating that noisy demonstrations impair the model’s ability to compress and extract task-relevant information. In the extreme case of 100% noise—where inputs and labels are completely uncorrelated—the model receives no meaningful task signal, and the characteristic compression-to-expression pattern disappears across layers. ($ii$) When the noise ratio remains below 40%, the minimum TDNV values stay below 1, indicating that within-task variance is still smaller than between-task distance. This allows task representations to remain distinguishable, resulting in minimal performance degradation. This observation explains the robust performance at noise ratios below 40%. However, beyond 40% noise, task representations become increasingly entangled, causing performance to decline rapidly.
The position of noisy demonstration affects compression.
@@
Unlike conventional machine learning, the order of demonstrations has a significant impact on model performance in ICL [43, 44, 45]. As illustrated in Figure˜11, perturbing demonstrations that appear later in the sequence causes a larger performance drop and higher TDNV values. These perturbations result in less compressed task vectors, indicating that later demonstrations play a more crucial role than earlier ones in helping the model extract task information. To gain deeper insight, we perform a fine-grained analysis by computing layerwise TDNV for each separator token, referred to as grid TDNV in Section˜D.2.
Figure 9: Layerwise TDNV and ICL accuracy under different perturbation positions.
Figure 10: Layerwise TDNV under different number of demonstrations $K$.
Figure 11: PCA of task vectors as number of demonstrations $K$ increase.